OpenStudy (anonymous):

Simplify the expression: 6x^2+7x-3 / 4x^2-9

OpenStudy (anonymous):

(2x+3)(3x-1)/(2x+3)(2x-3) = (3x-1)/(2x-3)

OpenStudy (luigi0210):

\[\frac{ 6x^2+7x-3 }{ 4x^2-9 }\] Factor out common multiples \[\frac{ (3x-1)(2x+3) }{ (2x+3)(2x-3) }\] Then simplify

OpenStudy (luigi0210):

So the final answer should be \[\frac{ 3x-1 }{ 2x-3 }\]

OpenStudy (anonymous):

and thats what i meant sorry

OpenStudy (anonymous):

Oh wow. Thanks Ya'll!!!

OpenStudy (luigi0210):

Same as physic said just in equation form :P